OpenStudy (christos):

For the functions f(x) = sqrt(x+1) and g(x)= 1/ sqrt(1-2x), find the domain of the function f/g

OpenStudy (anonymous):

\[\dfrac{\sqrt{x+1}\times\sqrt{1-2x}}{1}\] am i understanding you right?

OpenStudy (christos):

First time I am seeing this form :S

OpenStudy (anonymous):

I think you could write it on this form, then: find the possible x values

OpenStudy (christos):

what is this form equal to?

OpenStudy (anonymous):

x>-1 and x<1/2

OpenStudy (christos):

Would it be possible for you to help me continue? :/